# SegviGen: Repurposing 3D Generative Model for Part Segmentation

### 17 Mar 2026 v1

@@ 4.1 Setting
Figure 4: Full segmentation results. We compare SegviGen against a broad set of prior methods, where different colors indicate different segmented parts. From the results, SegviGen achieves high-accuracy full segmentation with sharp part boundaries using only 3D input. When 2D guidance is provided, it further allows explicit control over granularity and labels, enabling controllable, ultra-fine-grained 3D part parsing.
@@
Baselines. We compared our model’s performance on full segmentation between P3-SAM [43], Find3D [44], SAMPart3D [79], Partfield [38]. P3-SAM is a native 3D point-promptable part segmenter with multiple mask heads and an IoU predictor. It can be run automatically by sampling prompt points and merging redundant masks with NMS. Find3D targets open-world, language-queryable parts by auto-labeling rendered multi-view images with SAM and a VLM, projecting them back to 3D, and training a transformer to produce per-point features aligned to a CLIP-like embedding space for cosine-similarity querying. SAMPart3D and PartField both learn part-aware 3D features from multi-view SAM masks and obtain parts via feature clustering.
For interactive part segmentation, we compared our model against P3-SAM [43] and Point-SAM [86], where Point-SAM adapts the SAM prompt-and-mask paradigm to point clouds and is trained with SAM-generated pseudo masks.
@@ 4.2.2 Full Segmentation
We evaluate the full segmentation capability of SegviGen in two distinct settings: (1) Using purely native 3D representation. (2) Incorporating with 2D guidance. Quantitative comparisons with state-of-the-art methods, including Find3D [44], SAMPart3D [79], PartField [38], and P3-SAM [43], are presented in Table 2. Qualitative results are shown in 4.
@@ 5 Conclusion
Figure 7: Interactive demo. Users specify clicks on the 3D asset to perform interactive part segmentation, and can adjust visualization settings.

### 16 Apr 2026 v2

@@ 2 Related Work
There are also related works that focus on generating compositional and complete 3D parts. For example, HoloPart [80] assumes segmented parts as input for part completion, Dual Volume Packing [57] directly generates complete and separable parts from a single image, and BANG [86] recovers parts by first producing an exploded mesh sequence and then applying connected-component analysis and trajectory tracking. In contrast, SegviGen focuses on part segmentation, which often lays the foundation for such downstream tasks.
@@ 4.1 Setting
Figure 4: Full segmentation results. We compare SegviGen against a broad set of prior methods, where different colors indicate different segmented parts. From the results, SegviGen achieves high-accuracy full segmentation with sharp part boundaries using only 3D input. With 2D guidance, it further enables explicit control over granularity and labels for ultra-fine-grained 3D part parsing.
@@
Baselines. For full segmentation, we compare our method with P3-SAM [43], Find3D [44], SAMPart3D [82], and PartField [38]. P3-SAM is a native 3D point-promptable part segmenter with multiple mask heads and an IoU predictor. It can be run automatically by sampling prompt points and merging redundant masks with NMS. Find3D targets open-world, language-queryable parts by auto-labeling rendered multi-view images with SAM and a VLM, projecting them back to 3D, and training a transformer to produce per-point features aligned to a CLIP-like embedding space for cosine-similarity querying. SAMPart3D and PartField both learn part-aware 3D features from multi-view SAM masks and obtain parts via feature clustering. For interactive part segmentation, we compare our model with P3-SAM [43] and Point-SAM [90], the latter of which adapts the SAM prompt-and-mask paradigm to point clouds and is trained with SAM-generated pseudo masks.
@@ 4.2.2 Full Segmentation
We evaluate the full segmentation capability of SegviGen under two settings: (1) using only native 3D representations, and (2) incorporating 2D guidance. Quantitative comparisons with state-of-the-art methods are reported in Table 3, while qualitative results are shown in Fig. 4.
@@
Generalization on AI-Generated Meshes. Compared with manually modeled meshes, AI-generated meshes typically exhibit less well-defined part structures and weaker part separability, making part segmentation more challenging. To further evaluate the generalization ability of our method, we conduct experiments on synthetic test meshes generated by Hunyuan3D 2.1 [59]. Since ground-truth annotations are unavailable for these meshes, we perform a user study on 15 models with 47 experienced 3D practitioners, comparing our method against P3-SAM [43] in terms of segmentation plausibility and boundary accuracy. The support rates are reported in Table 2.
Table 2: User study on AI-generated meshes. “w. 2D Map” denotes the setting with 2D segmentation-map guidance.
[…]
P3-SAM [43] 15.7% 26.5%
SegviGen 29.0% 31.4%
SegviGen (w. 2D Map) 55.3% 42.1%
@@ 5 Conclusion
Figure 7: Interactive demo. Users specify clicks on the 3D asset to perform interactive part segmentation.
@@
6 Limitation and Future Work
Although SegviGen achieves strong overall performance, it still faces challenges when 2D guidance is unavailable. This mainly arises from the semantic–instance ambiguity in part segmentation, which can make it difficult for the model to produce sufficiently clear and satisfactory part boundaries. For future work, we believe that SegviGen can be naturally extended to other dense 3D prediction tasks by using color as a general representation of point-wise attributes. For instance, color could be used to encode stress magnitude for stress map prediction, or to represent per-vertex parameters associated with different bones, such as skinning weights, for skinning-related tasks. This indicates that the proposed paradigm may be applicable beyond segmentation to a broader class of dense 3D prediction problems.